OpenStudy (anonymous):

How do you calculate square roots of complex numbers? For example sqrt(-16-30i)

OpenStudy (epikrebel):

You want to square root -16-30i?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I checked on wolfram alpha and it told me the answer is 3-5i, but I don't know how to work it out.

OpenStudy (epikrebel):

3-5i?

OpenStudy (anonymous):

yes

hartnn (hartnn):

can u convert -16-30i into polar form?

hartnn (hartnn):

x+iy = -16-30i<--cartesian form r,theta<--polar form can u convert cartesian to polar form ?

OpenStudy (anonymous):

working on it

OpenStudy (anonymous):

-16-30i=(34,4.22)

hartnn (hartnn):

what is 4.22? in radians?

OpenStudy (anonymous):

yes

hartnn (hartnn):

ok, now to get square root of complex number, take square root of magnitude and divide the angle by 2

hartnn (hartnn):

that is here, take square root of 34

OpenStudy (anonymous):

ok, thanks

hartnn (hartnn):

so did u get 3-5i finally ?

OpenStudy (anonymous):

What is the reason for dividing the angle by 2? Must be a trig identity, right?

OpenStudy (anonymous):

I did! Thanks for your help.